What volume of alcohol with the density of 8.4×10²kg/m³ will have the same mass as 4.2m³ of petrol whose density is 7.2×10²kg/m³
1
Expert's answer
2019-05-30T09:10:32-0400

The mass of substance is related with density and volume as follows

"m=\rho V"

So


"\rho_1V_1=\rho_2V_2"

"V_2=V_1\frac{\rho_1}{\rho_2}"

"=4.2\:\rm{m^3}\times\frac{7.2\times 10^2\:\rm{kg/m^3}}{8.4\times 10^2\:\rm{kg/m^3}}"

"=3.6\:\rm{m^3}"